# Perpetual Future Contracts in Centralized and Decentralized Exchanges: Mechanism and Traders’ Behavior

### v2

@@ 5.3 Empirical Results and Discussion
Overall, an increase in unexpected liquidation in both long and short positions is expected to elevate price volatility. This is empirically corroborated in Table 4, where the estimated coefficient for unexpected liquidated volume on long positions is significantly positive ($5.008\times 10^{-7}$). Conversely, the aggregated effect of an increase in expected liquidation on short positions should theoretically be positive. However, this is contradicted by the negative sign of the estimated coefficient for expected liquidated volume on short positions ($-5.983\times 10^{-7}$). A plausible explanation for this anomaly is that a price rise, typically accompanying an increase in short position liquidations, may attract additional liquidity providers. This influx of liquidity providers potentially mitigates price volatility, with their dampening effect on volatility surpassing the amplifying impact of liquidations. We have conducted the Granger Causality Test to examine the relationship between between the return on Bitcoin and the net change in liquidity in the vBTC-vUSDC pool on Perpetual Protocol V2, which you can find in Appendix. It is shown that return on Bitcoin Granger-cause the net change in liquidity in the vBTC-vUSDC pool, while the reversed relationship does not hold.
@@
The asymmetrical effect of open interest on long and short positions is also predicted in the theoretical work by [43], which suggests that consuming liquidity involves a larger price impact than adding it, resulting in the larger marginal cost of a ETH buy order than the marginal benefit of a ETH sell order. The model proposed in [43] implies an ambiguous reaction of AMM liquidity to asset volatility, driven by the migration of both traders and liquidity providers between CEXs and DEXs. In our test, since liquidation explicitly includes behavior of both traders and liquidity providers, the ambiguity mentioned in [43] may help explain the inconsistency between our predicted effect of expected liquidation on short positions and the estimated coefficient.
5.3.4 Empirical results and analysis for GMX and GNS (with Oracle Pricing Model).
@@
In the realm of centralized exchanges (CEXs) and Perpetual Protocol V2, traders’ behaviors and informational inputs actively shape market prices. Conversely, in decentralized exchanges (DEXs) such as GMX and GNS, traders accept prices as a given, lacking the capacity to exert direct influence on price movements. The pricing mechanism in these DEXs is governed by a Price Oracle, as elaborated in Section 3.6, which assimilates prices from various exchanges, thereby rendering transactions on GMX and GNS incapable of impacting the price directly [^31] The code for GMX’s price feeding smart contract is publicly accessible at: https://github.com/gmx-io/gmx-contracts/blob/master/contracts/oracle/FastPriceFeed.sol.. As a result, trading activities on these platforms, encompassing trading volume and open interest, merely reflect traders’ interpretations and reactions to the information encapsulated within the Price Oracle’s determined prices.
@@
Drawing upon Shalen’s dispersion of beliefs model (1993) [2], which posits a market dichotomy between asymmetrically informed traders, market participants can be categorized into informed and uninformed traders [^32] Private information refers to the advantageous knowledge held predominantly by major market players such as fund companies and institutions, who have superior information gathering capabilities and insights into customer positions, distinguishing them from smaller retail investors.. Uninformed traders, lacking access to private information, often engage in irrational trading based on market noise and tend to overreact to information. They strive, albeit unsuccessfully, to discern private information and market trends from current price fluctuations, leading to heightened position-taking during periods of increased price volatility. In contrast, informed traders, operating on the basis of their private information, exhibit relatively consistent beliefs over time and typically engage in trading within a constrained price range, resulting in a negative correlation between their positions and price volatility [34].
@@
In Panel a of Table 4, the estimated coefficients that measure the relationship between open interest and price volatility are predominantly negative, except the case of long open interests in GNS. This indicates a more pronounced influence of informed traders’ behavior, leading to the negative signs in these estimations. The impact of uninformed traders on open interest is effectively counterbalanced by the actions of informed traders. Notably, traders’ behaviors exhibit asymmetry between buyers and sellers, as evidenced by less negative (in the case of GMX) or even insignificant (for GNS) coefficients for long positions. This asymmetry could be attributed to uninformed traders’ tendency to overreact more to positive news than negative, leading to an increased propensity for long position accumulation [^33] Existing literature on traditional spot and future markets provides evidence of traders’ asymmetric reactions to positive and negative news [44, 34]..
The asymmetry in long and short positions is further evidenced by variations in liquidated volume in relation to price volatility. As indicated in Table 4, for GMX and GNS, the coefficients linking liquidated volume on long positions are significantly positive, whereas those for short positions are not statistically significant. Additionally, Panel b of Table 4 reveals that long positions are more likely to liquidation than short positions, as denoted by the significantly positive coefficients for $\mathcal{LIQ}_{\text{long }}$ compared to the insignificant coefficients for $\mathcal{LIQ}_{\text{short }}$. This suggests that liquidation on long positions escalates both in frequency and magnitude when price volatility intensifies, corroborating the notion of traders’ asymmetric overreaction to good news. Traders tend to overreact by not only increasing long positions but also by employing higher leverage for long positions compared to short, as evidenced in Table 3, resulting in a higher rate of liquidation for long positions in both volume and probability.
@@
Aligned with Shalen’s model, it is anticipated that uninformed traders are more likely to increase leverage during periods of heightened price volatility to maximize capital efficiency, thereby enlarging their positions at a lower cost. In contrast, informed traders are less inclined to risk capital under volatile conditions, leading to a negative correlation between leverage and price volatility. Table 5 delineates the estimated results for Eq.(31) and Eq.(32). All eight coefficients relating price volatility and average leverage are estimated to be negative (as shown in the two columns under Regression on Eq.(31)), indicating a more pronounced change in leverage as influenced by informed traders compared to uninformed traders. This inference is further substantiated by the estimation results for Eq.(32), which demonstrate that trading volumes with lower leverages (less than 50×50$\times$50 ×) increase, while those with higher leverages (more than 50×50$\times$50 ×) remain statistically unchanged in response to rising price volatility.
In summation, the analysis reveals that informed traders exhibit a higher level of activity compared to their uninformed counterparts on GMX and GNS, a conclusion substantiated by empirical data pertaining to the dynamics of open interests and leverages. Additionally, the observed asymmetry in open interests and liquidations between long and short positions suggests a tendency among traders to disproportionately overreact to good news. However, it is imperative to acknowledge a critical distinction for analytical purposes: unlike GMX and GNS, centralized exchanges (CEXs) possess a price discovery function, leading to a reciprocal influence between traders’ behavior and market prices. Under Efficient-Market Hypothesis [45], the price aggregate and reflect all the current private information and information contained in the past pricing history. Consequently, it is methodologically unsound to interpret changes in trading activities, e.g., volume, open interest, liquidation, and leverage, solely as reactions to price fluctuations. This perspective challenges the conclusions drawn by [21], who posited that changes in liquidation are merely passive responses to price movements, thereby overlooking the role of liquidation as a contributory factor in price formation within BitMEX, a CEX. This oversight underscores the necessity for a nuanced understanding of the interplay between trader behavior and price dynamics, particularly in the context of CEXs where the price discovery process is inherently more complex.
Table 5: Regression results on Eq. (31) and Eq. (32)
Perpetual Contracts on Bitcoin
Regression on Eq. (31) Regression on Eq. (32)
Variables GMX GNS GMX GNS
Intercept 2.287e-12 (9.47)*** 0.028 (3.84)*** 1.609e-12 (15.55)*** 9.729e-12 (3.89)***
Lagged volatility 0.229 (7.26)*** 0.369 (10.66)*** 0.200 (6.37)*** 0.374 (10.05)***
@@
Expected trading volume 1.051e-10 (5.60)*** 5.201e-11 (0.83) - -
Unexpected trading volume 1.359e-10 (10.41)*** 4.593e-10 (6.86)*** - -
Expected OI on short -4.405e-10 (-6.52)*** -1.626e-10 (-1.58) -5.415e-10 (-7.32)*** -1.893e-10 (-1.67)*
Unexpected OI on short -2.485e-10 (-2.89)*** -4.423e-10 (-2.98)*** -2.767e-10 (-3.01)*** -4.355e-10 (-2.71)***
Expected OI on long -1.373e-10 (-4.11)*** 3.964e-11 (0.58) -2.318e-10 (-4.76)*** 2.426e-11 (0.33)
Unexpected OI on long -1.091e-10 (-1.36) -5.505e-11 (-0.38) -2.97e-10 (-3.61)*** 4.033e-11 (0.26)
@@
Expected liquidated volume on short positions 2.538e-09 (1.15) -1.467e-09 (-1.27) 4.888e-10 (0.21) 5.541e-10 (0.40)
Unexpected liquidated volume on short positions 1.891e-09 (1.55) 3.177e-10 (0.80) 2.289e-09 (1.86)* 1.925e-10 (0.42)
Expected liquidated volume on long positions 2.77e-07 (9.14)*** 7.573e-09 (3.45)*** 1.965e-07 (14.98)*** 5.472e-09 (2.27)**
Unexpected liquidated volume on long positions 2.539e-09 (5.06)*** 1.384e-09 (1.42) 2.455e-09 (4.78)*** 3.984e-09 (3.70)***
@@
Expected average leverage on short positions -0.0004 (-2.21)** -0.0001 (-1.40) - -
Unexpected average leverage on short positions -0.0002 (-3.00)*** -0.0002 (-4.37)*** - -
Expected average leverage on long positions -0.0004 (-2.15)** -1.193e-05 (-0.36) - -
Unexpected average leverage on long positions -0.0003 (-3.65)*** -7.237e-05 (-1.93)** - -
Leverage (in categories):
Expected: less than 1×1$\times$1 × - - -2.112e-08 (-1.21) -
Unexpected: less than 1×1$\times$1 × - - 1.688e-08 (3.46)*** -
Expected: 1×1$\times$1 × to 5×5$\times$5 × - - 1.829e-10 (1.08) 1.684e-08 (0.23)
Unexpected: 1×1$\times$1 × to 5×5$\times$5 × - - 9.772e-10 (5.51)*** -1.289e-08 (-0.54)
Expected: 5×5$\times$5 × to 10×10$\times$10 × - - 1.847e-09 (3.44)*** -2.534e-08 (-0.55)
Unexpected: 5×5$\times$5 × to 10×10$\times$10 × - - 1.012e-09 (5.27)*** 1.736e-08 (1.08)
Expected: 10×10$\times$10 × to 15×15$\times$15 × - - -6.697e-10 (-1.50) 5.39e-08 (0.96)
Unexpected: 10×10$\times$10 × to 15×15$\times$15 × - - 8.724e-11 (0.63) 6.408e-09 (0.44)
Expected: 15×15$\times$15 × to 20×20$\times$20 × - - 1.998e-09 (3.75)*** 9.314e-08 (1.84)*
Unexpected: 15×15$\times$15 × to 20×20$\times$20 × - - 2.023e-10 (1.28) 5.102e-09 (0.33)
Expected: 20×20$\times$20 × to 25×25$\times$25 × - - -1.06e-10 (-0.12) 2.698e-07 (3.89)***
Unexpected: 20×20$\times$20 × to 25×25$\times$25 × - - 2.66e-10 (2.15)** 6.261e-09 (0.92)
Expected: 25×25$\times$25 × to 30×30$\times$30 × - - -9.034e-11 (-0.76) -4.597e-08 (-1.31)
Unexpected: 25×25$\times$25 × to 30×30$\times$30 × - - 9.099e-11 (1.70)** -7.329e-09 (-0.91)
Expected: 30×30$\times$30 × to 35×35$\times$35 × - - 5.761e-10 (1.91)* 8.934e-08 (1.34)
Unexpected: 30×30$\times$30 × to 35×35$\times$35 × - - -3.618e-12 (-0.02) -3.778e-09 (-0.26)
Expected: 35×35$\times$35 × to 40×40$\times$40 × - - 2.882e-11 (0.03) 2.465e-07 (3.10)***
Unexpected: 35×35$\times$35 × to 40×40$\times$40 × - - 2.266e-10 (0.93) -1.913e-09 (-0.32)
Expected: 40×40$\times$40 × to 45×45$\times$45 × - - 1.289e-09 (0.51) -2.922e-10 (-0.01)
Unexpected: 40×40$\times$40 × to 45×45$\times$45 × - - 3.24e-10 (1.12) -2.72e-09 (-0.61)
Expected: 45×45$\times$45 × to 50×50$\times$50 × - - 2.626e-11 (0.15) 1.235e-08 (0.74)
Unexpected: 45×45$\times$45 × to 50×50$\times$50 × - - 1.692e-10 (1.73)* 9.194e-10 (0.30)
Expected: more than 50×50$\times$50 × - - -1.529e-10 (-0.29) -1.466e-09 (-0.13)
Unexpected: more than 50×50$\times$50 × - - -3.74e-11 (-0.20) -1.21e-08 (-0.96)
Adjusted $R^{2}$ 0.398 0.339 0.427 0.260
AIC -4405 -3693 -4424 -3608
No. of obs. 747 628 747 628
Notes. This table reports regression results on Eq. (31) and Eq. (32). For the regression with Eq. (32), Leverage (in categories) means trading volumes in leverage within different ranges, which are grouped into categories. For GNS, our dataset does not include any transaction with leverage less than 1×1$\times$1 ×, so we drop the variables of trading volumes with the leverage less than 1×1$\times$1 ×. In each cell, the t-statistics are in the parentheses. ***, ***** *, and ******* * * denote significance at 0.1, 0.05, and 0.01 level, respectively.
@@ 6 Conclusion
In conclusion, this study presents an in-depth Systematic Survey of Knowledge (SoK) that effectively contrasts the operational dynamics of Centralized Exchanges (CEXs) and Decentralized Exchanges (DEXs) in the realm of perpetual futures trading. It significantly enriches the academic discourse by meticulously delineating the design elements, order matching mechanisms, price discovery processes, and custody of funds in both CEXs and DEXs, thereby offering a robust foundation for future research in this field. With the models in our framework, any exchange could be classified into these predined models and each model corresponds to a sort of distinctive features. The study underscores the necessity of developing new theoretical models, moving beyond conventional frameworks like Kyle’s model[1], to aptly comprehend the intricacies of DEX models. Through an empirical analysis of transaction data from August 2020 to September 2023, it reveals a positive correlation between price volatility and trading volumes in exchanges of CEX Model, in line with traditional futures markets. However, it also uncovers distinct variations in DEXs, particularly under the VAMM Pricing Model, where open interest impacts long and short positions differently. The behavior of traders in GMX and GNS, who operate under Oracle Pricing and act as price takers, is found to be consistent with Shalen’s dispersion of beliefs model[2]. Together with Table 2, Table LABEL:tab:summary_empirical_results summarizes traders’ distinctive behavior derived by different design of the models. Furthermore, the study highlights the propensity of uninformed traders to disproportionately react to positive news than negative news, thereby deepening our understanding of trader behavior in DEXs and offering valuable insights for the development of future theoretical models in this evolving domain.
@@
Exchange Model Exchange Analyzed Price Discovering Price Volatility & Trading Volume Price Volatility & Open Interest Symmetry; Explanation
CEX Model Binance; Bybit Yes Positively correlated Negatively correlated Symmetrical; Kyle’s model [1]
Hybrid Model - Yes - - -
VAMM Model Perpetual Prorocol V2 Yes Positively correlated
Negatively correlated for short positions & positively correlated for long positions
(in the long-run); positive (in the short-run)
Asymmetrical; Uniswap v3 AMM
Oracle Pricing Model GMX; GNS Yes Positively correlated Negatively correlated
Asymmetrical; Shalen’s dispersion of beliefs model [2]
In terms of future work, firstly, the observed asymmetry in trading behaviors between buyers and sellers within the Oracle Price Model DEX warrants a deeper exploration into the psychological underpinnings of this phenomenon. Future research could delve into whether uninformed traders disproportionately react to positive news compared to negative news, and examine the extent to which cognitive biases or prevailing social norms may influence such behaviors. This line of inquiry could provide valuable insights into the psychological mechanisms at play in decentralized exchange environments.
Secondly, the role of liquidity providers, a novel concept in the decentralized finance (DeFi) ecosystem, presents a rich area for academic investigation. While this paper primarily analyzes the behaviors of traders in Perpetual Future Contracts, a comparative study focusing on their counterparts, the liquidity providers, could yield significant findings. Future research should consider the impact of variables such as interest rates and the availability of alternative DeFi products on the behaviors of liquidity providers, particularly in the context of Perpetual Future Contracts. Understanding how these factors interact to influence liquidity provider behaviors could offer a more comprehensive view of market dynamics in DeFi.
Lastly, the potential of blockchain-based financial instruments to enhance transparency and accountability in micro financial markets is an important avenue for future research. The decentralized and transparent nature of blockchain technology suggests that these instruments could be instrumental in promoting accountability and transparency in micro financial markets, potentially encouraging more ethical investment practices. Future studies should explore the capacity of blockchain-based instruments to achieve these outcomes, examining the factors that may affect their effectiveness in fostering responsible financial practices.

### v4

@@ 3.1 Empirical results and analysis for LOB Model (Binance and Bybit)
Overall, an increase in unexpected liquidation in both long and short positions is expected to elevate price volatility. This is empirically corroborated in Table 1, where the estimated coefficient for unexpected liquidated volume on long positions is significantly positive ($5.008\times 10^{-7}$). Conversely, the aggregated effect of an increase in expected liquidation on short positions should theoretically be positive. However, this is contradicted by the negative sign of the estimated coefficient for expected liquidated volume on short positions ($-5.983\times 10^{-7}$). A plausible explanation for this anomaly is that a price rise, typically accompanying an increase in short position liquidations, may attract additional liquidity providers. This influx of liquidity providers potentially mitigates price volatility, with their dampening effect on volatility surpassing the amplifying impact of liquidations. We have conducted the Granger Causality Test to examine the relationship between between the return on Bitcoin and the net change in liquidity in the vBTC-vUSDC pool on Perpetual Protocol V2, which you can find in Appendix B. It is shown that return on Bitcoin Granger-cause the net change in liquidity in the vBTC-vUSDC pool, while the reversed relationship does not hold.
@@
The asymmetrical effect of open interest on long and short positions is also predicted in the theoretical work by [3], which suggests that consuming liquidity involves a larger price impact than adding it, resulting in the larger marginal cost of a ETH buy order than the marginal benefit of a ETH sell order. The model proposed in [3] implies an ambiguous reaction of AMM liquidity to asset price volatility, driven by the migration of both traders and liquidity providers between CEXs and DEXs. In our test, since liquidation explicitly includes behavior of both traders and liquidity providers, the ambiguity mentioned in [3] may help explain the inconsistency between our predicted effect of expected liquidation on short positions and the estimated coefficient.
3.3 Empirical results and analysis for Oracle Pricing Model (GMX and GNS).
@@
In exchanges of Oracle Pricing Model such as GMX and GNS, traders accept prices as a given, lacking the capacity to exert direct influence on price movements. The pricing mechanism in these DEXs is governed by a Price Oracle, which assimilates prices from various exchanges, thereby rendering transactions on GMX and GNS incapable of impacting the price directly [^13] The code for GMX’s price feeding smart contract is publicly accessible at: https://github.com/gmx-io/gmx-contracts/blob/master/contracts/oracle/FastPriceFeed.sol.. As a result, trading activities on these platforms, encompassing trading volume and open interest, merely reflect traders’ interpretations and reactions to the information encapsulated within the Price Oracle’s determined prices.
@@
Drawing upon Shalen’s dispersion of beliefs model (1993) [18], which posits a market dichotomy between asymmetrically informed traders, market participants can be categorized into informed and uninformed traders. Uninformed traders, lacking access to private information, often engage in irrational trading based on market noise and tend to overreact to information. They strive, albeit unsuccessfully, to discern private information and market trends from current price fluctuations, leading to heightened position-taking during periods of increased price volatility. In contrast, informed traders, operating on the basis of their private information, exhibit relatively consistent beliefs over time and typically engage in trading within a constrained price range, resulting in a negative correlation between their positions and price volatility [24].
@@
In Table 1, the estimated coefficients that measure the relationship between open interest and price volatility are predominantly negative, except the case of long open interests in GNS. This indicates a more pronounced influence of informed traders’ behavior, leading to the negative signs in these estimations. The impact of uninformed traders on open interest is effectively counterbalanced by the actions of informed traders. Notably, traders’ behaviors exhibit asymmetry between buyers and sellers, as evidenced by less negative (in the case of GMX) or even insignificant (for GNS) coefficients for long positions. This asymmetry could be attributed to uninformed traders’ tendency to overreact more to positive news than negative, leading to an increased propensity for long position accumulation [^14] Existing literature on traditional spot and future markets provides evidence of traders’ asymmetric reactions to positive and negative news [15, 24].. This asymmetry is further evidenced by variations in liquidated volume in relation to price volatility. As indicated in Table 1, for GMX and GNS, the coefficients linking liquidated volume on long positions are significantly positive, whereas those for short positions are not statistically significant.
@@
Aligned with Shalen’s model, it is anticipated that uninformed traders are more likely to increase leverage during periods of heightened price volatility to maximize capital efficiency, thereby enlarging their positions at a lower cost. In contrast, informed traders are less inclined to risk capital under volatile conditions, leading to a negative correlation between leverage and price volatility. Table 2 delineates the estimated results for Eq.(3). All eight coefficients relating price volatility and average leverage are estimated to be negative, indicating a more pronounced change in leverage as influenced by informed traders compared to uninformed traders.
Table 2: Regression results on Eq. (3)
[…]
Intercept [^17] In each cell, the t-statistics are in the parentheses. *, **∗ ∗, and ***∗ ∗ ∗ denote significance at 0.1, 0.05, and 0.01 level, respectively. 2.287e-12 (9.47)*** 0.028 (3.84)*** Lagged volatility 0.229 (7.26)*** 0.369 (10.66)***
Expected portion Unexpected portion
@@
Trading volume 1.051e-10 (5.60)*** 5.201e-11 (0.83) 1.359e-10 (10.41)*** 4.593e-10 (6.86)***
OI on short -4.405e-10 (-6.52)*** -1.626e-10 (-1.58) -2.485e-10 (-2.89)*** -4.423e-10 (-2.98)***
OI on long -1.373e-10 (-4.11)*** 3.964e-11 (0.58) -1.091e-10 (-1.36) -5.505e-11 (-0.38)
@@
Liquidated volume on short positions 2.538e-09 (1.15) -1.467e-09 (-1.27) 1.891e-09 (1.55) 3.177e-10 (0.80)
Liquidated volume on long positions 2.77e-07 (9.14)*** 7.573e-09 (3.45)*** 2.539e-09 (5.06)*** 1.384e-09 (1.42)
@@
Average leverage on short positions -0.0004 (-2.21)** -0.0001 (-1.40) -0.0002 (-3.00)*** -0.0002 (-4.37)***
Average leverage on long positions -0.0004 (-2.15)** -1.193e-05 (-0.36) -0.0003 (-3.65)*** -7.237e-05 (-1.93)**
Adjusted $R^{2}$ 0.398 0.339 AIC -4405 -3693
No. of obs. 747 628
In summary, the analysis reveals that informed traders exhibit a higher level of activity compared to their uninformed counterparts on GMX and GNS, a conclusion substantiated by empirical data pertaining to the dynamics of open interests and leverages. Additionally, the observed asymmetry in open interests and liquidations between long and short positions suggests a tendency among traders to disproportionately overreact to good news. However, it is imperative to acknowledge a critical distinction for analytical purposes: unlike GMX and GNS, centralized exchanges (CEXs) possess a price discovery function, leading to a reciprocal influence between traders’ behavior and market prices. Under Efficient-Market Hypothesis [23], the price aggregate and reflect all the current private information and information contained in the past pricing history. Consequently, it is methodologically unsound to interpret changes in trading activities, e.g., volume, open interest, liquidation, and leverage, solely as reactions to price fluctuations. This perspective challenges the conclusions drawn by [20], who posited that changes in liquidation are merely passive responses to price movements, thereby overlooking the role of liquidation as a contributory factor in price formation within BitMEX, a CEX. This oversight underscores the necessity for a nuanced understanding of the interplay between trader behavior and price dynamics, particularly in the context of CEXs where the price discovery process is inherently more complex.
@@ 4 Conclusion
This study underscores the necessity of developing new theoretical models, moving beyond conventional frameworks like Kyle’s model[13], to aptly comprehend the intricacies of DEXs. Our empirical results uncover distinct variations in DEXs, particularly under the VAMM Model, where open interest impacts long and short positions differently. The behavior of traders in exchanges of Oracle Pricing Model (GMX and GNS), who act as price takers, is found to be consistent with Shalen’s dispersion of beliefs model[18]. Our study highlights the significant fundamental difference between CEXs and DEXs, based on which theoretical frameworks specifically for DEX design can be derived in the future works. Table 3 encapsulates a summary of the unique trading behaviors elicited by the varied designs of these models. Any exchange classified into one of our three exchange models is believed to derive distinctive behavior of the traders caused by the design of the exchange model.
@@
[…]
Analyzed
LOB Model Yes Positively correlated Negatively correlated Symmetrical; Kyle’s model [13] Binance; Bybit
VAMM Model Yes Positively correlated Mixed Asymmetrical; Uniswap v3 AMM Perpetual Prorocol V2
Oracle Pricing Model No Positively correlated Negatively correlated Asymmetrical; Shalen’s dispersion of beliefs model [18] GMX; GNS